1. 3-4q=10q+10
2. 3k-5=7k-21
3. 5t-9=3t+7
4. 8s+9=7s+6
its solving questions on each side of the problem

Answer by piglet162431(90)

1. 3-4q=10q+10
14q=-7
q=-1/2
2. 3k-5=7k-21
-4k=-16
k=4
3. 5t-9=3t+7
2t=16
t=8
4. 8s+9=7s+6
s=-3
For specification plz email. I just answered this in a hurry.